# Automated event generation for S-wave quarkonium and leptonium production in NRQCD and NRQED

### v1

@@ 7.2 True muonium production
Table 14: Cross-section predictions for para- ($\mathit{TM}_{0}$) and ortho-true muonium ($\mathit{TM}_{1}$) production at $\sqrt{s}=13$ TeV in $pp$ collisions and at $\sqrt{s}=10.58\,\mathrm{GeV}$ in $e^{+}e^{-}$ collisions, based on the setup described in section 4. For ortho-true muonium, additional fiducial cuts are applied: $p_{T,j}>2\,\mathrm{GeV}$ and $|\eta_{j}|<5$ for final-state jets, and $p_{T,\gamma}>0.5\,\mathrm{GeV}$ and $|\eta_{\gamma}|<5$ for photons. All cross sections are rescaled to the $\alpha(0)$ scheme. Numbers in parentheses indicate numerical uncertainties from the MC phase-space integration.
Following our analysis of positronium, we present predictions for the ground-state configurations of $n={}^{1}S_{0}$ para-true muonium ($\mathit{TM}_{0}$) and $n={}^{3}S_{1}$ ortho-true muonium ($\mathit{TM}_{1}$). In table 14, we show the cross section for the hadronic process
@@ 7.3 True tauonium production
As in the cases of positronium and true muonium discussed in the previous sections, we present cross-section predictions for ditauonium production in its ground-state configurations with principal quantum number $N=1$. Specifically, we consider the $n={}^{1}S_{0}$ singlet para-ditauonium, denoted $\mathcal{T}_{0}$, and the $n={}^{3}S_{1}$ triplet ortho-ditauonium, $\mathcal{T}_{1}$. Calculations are performed using the baseline setup described in section 4 for both $e^{+}e^{-}$ and $pp$ colliders. A summary of the results is provided in table 15.
@@
Table 15: Cross-section predictions for para- ($\mathcal{T}_{0}$) and ortho-ditauonium ($\mathcal{T}_{1}$) production at $\sqrt{s}=13$ TeV in $pp$ collisions and at $\sqrt{s}=10.58$ GeV in $e^{+}e^{-}$ collisions, based on the setup described in section 4. Additional fiducial cuts, $p_{T,j}>2\,\mathrm{GeV}$ and $|\eta_{j}|<5$ ($p_{T,\gamma}>0.5\,\mathrm{GeV}$ and $|\eta_{\gamma}|<5$), are applied to final-state jets (photons) in the computations for ortho-ditauonium production. All cross sections are rescaled to the $\alpha(0)$ scheme. Numbers in parentheses indicate numerical uncertainties from the MC phase-space integration.
@@ 8 Conclusions
Benchmark comparisons with existing tools and analytic expressions confirm the accuracy and reliability of our implementation. Cross-section predictions have been provided for a broad set of final-state bound states and observables, demonstrating the potential of our extension. In addition, our implementation is fully compatible with and seamlessly integrates into other MG5_aMC features, including custom UFO models adapted for bound-state production, the parton-shower interface, and gamma-UPC for studying photon–photon collisions in proton or nuclear UPCs.
Our work marks an important first step toward delivering state-of-the-art theoretical predictions for bound-state production studies within the collinear factorisation framework in the widely-used event generator MG5_aMC. Future studies include the automation of P-wave quarkonium production and, in the near term, extension to NLO accuracy. These developments will further expand the scope of bound-state studies in collider environments and facilitate future global NRQCD analyses and comparisons with experimental data. All implementations will be made publicly available via the standard MG5_aMC distribution on Launchpad  [^32] https://launchpad.net/mg5amcnlo, as well as via the NLOAccess EU Virtual Access  [^33] https://nloaccess.in2p3.fr/, which provides the high-energy, hadronic, and heavy-ion physics communities with versatile, high-precision tools for quarkonium and leptonium physics in a user-accessible format.

### v2

@@ 7.2 True muonium production
Table 14: Cross-section results for para- ($\mathit{TM}_{0}$) and ortho-true muonium ($\mathit{TM}_{1}$) production at $\sqrt{s}=13$ TeV in $pp$ collisions and at $\sqrt{s}=10.58\,\mathrm{GeV}$ in $e^{+}e^{-}$ collisions, based on the setup described in section 4. For ortho-true muonium, additional fiducial cuts are applied: $p_{T,j}>2\,\mathrm{GeV}$ and $|\eta_{j}|<5$ for final-state jets, and $p_{T,\gamma}>0.5\,\mathrm{GeV}$ and $|\eta_{\gamma}|<5$ for photons. All cross sections are rescaled to the $\alpha(0)$ scheme. Numbers in parentheses indicate numerical uncertainties from the MC phase-space integration.
Following our analysis of positronium, we present cross sections for the ground-state configurations of $n={}^{1}S_{0}$ para-true muonium ($\mathit{TM}_{0}$) and $n={}^{3}S_{1}$ ortho-true muonium ($\mathit{TM}_{1}$). In table 14, we show the cross section for the hadronic process
@@ 7.3 True tauonium production
As in the cases of positronium and true muonium discussed in the previous sections, we present cross-section calculations for ditauonium production in its ground-state configurations with principal quantum number $N=1$. Specifically, we consider the $n={}^{1}S_{0}$ singlet para-ditauonium, denoted $\mathcal{T}_{0}$, and the $n={}^{3}S_{1}$ triplet ortho-ditauonium, $\mathcal{T}_{1}$. Calculations are performed using the baseline setup described in section 4 for both $e^{+}e^{-}$ and $pp$ colliders. A summary of the results is provided in table 15.
@@
Table 15: Cross-section results for para- ($\mathcal{T}_{0}$) and ortho-ditauonium ($\mathcal{T}_{1}$) production at $\sqrt{s}=13$ TeV in $pp$ collisions and at $\sqrt{s}=10.58$ GeV in $e^{+}e^{-}$ collisions, based on the setup described in section 4. Additional fiducial cuts, $p_{T,j}>2\,\mathrm{GeV}$ and $|\eta_{j}|<5$ ($p_{T,\gamma}>0.5\,\mathrm{GeV}$ and $|\eta_{\gamma}|<5$), are applied to final-state jets (photons) in the computations for ortho-ditauonium production. All cross sections are rescaled to the $\alpha(0)$ scheme. Numbers in parentheses indicate numerical uncertainties from the MC phase-space integration.
@@ 8 Conclusions
Benchmark comparisons with existing tools and analytic expressions confirm the accuracy and reliability of our implementation. Cross-section computations have been provided for a broad set of final-state bound states and observables, demonstrating the potential of our extension. In addition, our implementation is fully compatible with and seamlessly integrates into other MG5_aMC features, including custom UFO models adapted for bound-state production, the parton-shower interface, and gamma-UPC for studying photon–photon collisions in proton or nuclear UPCs.
Our work represents an important first step toward delivering state-of-the-art theoretical predictions for bound-state production studies within the collinear factorisation framework in the widely-used event generator MG5_aMC. Future studies include the automation of P-wave quarkonium production and, in the near term, extension to NLO accuracy. These developments will further expand the scope of bound-state studies in collider environments and facilitate future global NRQCD analyses and comparisons with experimental data. All implementations will be made publicly available via the standard MG5_aMC distribution on Launchpad  [^32] https://launchpad.net/mg5amcnlo, as well as via the NLOAccess EU Virtual Access  [^33] https://nloaccess.in2p3.fr/, which provides the high-energy, hadronic, and heavy-ion physics communities with versatile, high-precision tools for quarkonium and leptonium physics in a user-accessible format.